### v2

# Bound of Casimir Effect by Holography

Abstract
Inspired by the Kovtun-Son-Starinet bound, we propose that holography impose a lower bound on the Casimir effect. For simplicity, we focus on the Casimir effect between parallel planes for 3d conformal field theories and briefly comment on the generalizations to other boundary shapes and higher dimensions. Remarkably, the ghost-free holographic models impose a universal lower bound of the Casimir effect. We verify the holographic bound by free theories, Ising model, and $O(N)$ model with $N=2,3$ at critical points. Remarkably, the holographic bound is also obeyed by a general class of quantum field theories without conformal symmetries. It is interesting to find a field-theoretical proof or counterexample for the holographic bound of Casimir effect.
Introduction — Casimir effect Casimir:1948dh is a novel quantum effect that originates from the changes of vacuum fluctuations due to the boundary Plunien:1986ca ; Bordag:2001qi ; Milton:2004ya ; Bordag:2009zz . It is a candidate of dark energy Wang:2016och and plays a vital role in QCD Vepstas:1984sw and worm-hole physics Morris:1988tu ; Maldacena:2020sxe . It has been measured in experiments Mohideen:1998iz ; Bressi:2002fr ; Klimchitskaya:2009cw and has significant potential applications in nanotechnology. As the vacuum energy, Casimir energy is defined as the lowest energy of a quantum system with a boundary. Since energy should be bounded from the below, it is expected that there is a fundamental lower bound of the Casimir effect. This letter explores this vital problem. Holography (AdS/CFT) Maldacena:1997re provides a powerful tool to study the strongly coupled conformal field theory (CFT) and usually imposes bounds on various physical phenomena. The most famous example is the Kovtun-Son-Starinet (KSS) bound Kovtun:2004de for hydromechanics, which conjectures that the ratio of shear viscosity $\eta$ to entropy density $s$ has a lower bound
$\frac{\eta}{s}\geq\frac{\hbar}{4\pi k_{B}},$ (1)
where $\hbar$ is the reduced Planck constant and $k_{B}$ is the Boltzmann constant. The lower bound is obtained from the strongly-coupled CFT dual to Einstein gravity Policastro:2001yc and can be modified slightly by higher derivative gravity Brigante:2007nu ; Kats:2007mq ; Brigante:2008gz .
Now, let us explore whether AdS/CFT set a similar lower bound for the Casimir effect. For simplicity, we focus on the Casimir effect for 3d CFTs between parallel planes (strip) and comment on various generalizations at the end of the letter. We consider flat space and impose the same boundary conditions on the two planes. Then, the Casimir effect takes a universal form
$\langle T^{i}_{\ j}\rangle_{\text{strip}}=\kappa_{1}\frac{\hbar c}{L^{3}}\text{diag}\Big{(}1,-2,1\Big{)},$ (2)
where $\kappa_{1}$ is the dimensionless Casimir amplitude, $c$ is the speed of light, $L$ is the distance between the two plates. For a CFT with fixed strip width $L$, $(-\kappa_{1})$ should be bounded from below to avoid negative infinity energy density $(T_{tt}=-\kappa_{1}\hbar c/L^{3})$ and pressure $(T_{nn}=-2\kappa_{1}\hbar c/L^{3})$. Note that the Casimir amplitude $\kappa_{1}$ is an extensive parameter in the sense that if one CFT has a Casimir amplitude $\kappa_{1}$, then $N$-copies of the same CFT has a Casimir amplitude $N\kappa_{1}$. It implies only $(-\kappa_{1})$ over another extensive parameter could have a lower bound. Natural candidates of these extensive parameters are the central charges of CFTs. Since the Casimir effect is a boundary effect, the most natural candidates are boundary central charges defined in Weyl anomalyJensen:2015swa $\mathcal{A}=\int_{\partial M}d^{2}x\sqrt{|\sigma|}\Big{(}aR_{\partial M}+b\text{tr}\bar{k}^{2}\Big{)},$ where $R_{\partial M}$ and $\bar{k}$ are the intrinsic Ricci scalar and traceless parts of extrinsic curvatures on the boundary $\partial M$, respectively. $a$ is the A-type boundary central charge, which decreases under the boundary RG flow Jensen:2015swa . We remark that $a$ can be positive, zero, or negative. For example, $a=\pm 1/(384\pi)$ for conformally coupled free scalar with Robin boundary condition (RBC) and Dirichlet boundary condition (DBC) Jensen:2015swa , respectively; $a\sim\sinh(\rho)$ with $-\infty<\rho<\infty$ for holographic CFT3 Fujita:2011fp . Since $a$ can be zero and negative, there is no lower bound for $(-\kappa_{1}/a)$. On the other hand, the B-type boundary central charge $b$ is always positive since it is related to the Zamolodchikov norm of displacement operator, i.e., $b=C_{D}\pi/32>0$. The displacement operator $D(y)$ describes the violation of space translation invariance normal to the boundary Billo:2016cpy , i.e., $\nabla_{i}T^{ij}=\delta(x)D(y)n^{j}$. The two-point function defines its Zamolodchikov norm, i.e., $\langle D(y)D(0)\rangle=C_{D}/|y|^{6}$. By definition, $C_{D}=32b/\pi$ is always positive. Thus, we propose the ratio of Casimir amplitude to norm of displacement operator has a lower bound set by holography
[…]
Remarkably, unlike the KSS bound (1), different holographic models give the same universal lower bound (3) for Casimir effect. We verify the lower bound (3) by free CFTs, Ising model, and $O(N)$ model with $N=2,3$ at critical points. See Table. 1. Below we discuss the derivation of the holographic lower bound and its various tests and generalizations. We focus on natural units with $c=\hbar=k_{B}=1$ in the followings.
Table 1: $(-\kappa_{1}/C_{D})$ for various 3d CFTs
Fermion Scalar Ising O(2) O(3) Holography
$-1.42$ $-1.89$2.12 -1.83 -1.62 $-2.17$
Holographic bound — Let us study the holographic bound of the Casimir effect of a strip in AdS/BCFT Takayanagi:2011zk . We are interested in the general ghost-free gravity in AdS4, which turns out to be Dvali-Gabadadze-Porrati (DGP) gravity Dvali:2000hr $I=\int_{N}d^{4}x\sqrt{|g|}\Big{(}R+6\Big{)}+2\int_{Q}d^{3}y\sqrt{|h|}\Big{(}K-T+\lambda\mathcal{R}\Big{)}$, where $K$ is the extrinsic curvature, $T$ is the brane tension, $\lambda$ is the DGP parameter, $R$ and $\mathcal{R}$ are Ricci scalars in bulk $N$ and on the brane $Q$ respectively. We have set Newton’s constant $116\pi G_{N}=$ 1 and AdS radius $l=1$ for simplicity. Note that higher derivative gravity and negative DGP gravity generally suffer the ghost problems Miao:2023mui , while Gauss-Bonnet/Lovelock gravity appears only in dimensions higher than four. Thus, positive DGP gravity with $\lambda\geq 0$ is the general ghost-free gravity theory in AdS4. Following Takayanagi:2011zk , we impose Neumann boundary condition (NBC) on the brane $Q$
$K^{ij}-(K-T+\lambda\mathcal{R})h^{ij}+2\lambda\mathcal{R}^{ij}=0,$ (4)
where $T=2\tanh(\rho)-2\lambda\text{sech}^{2}(\rho)$ is a useful parameterization Miao:2023mui . The holographic norm of the displacement operator is given by Miao:2023mui ; Miao:2018dvm
$C_{D}=\frac{32}{\pi\left(\frac{2\lambda}{2\lambda\sinh(\rho)+\cosh(\rho)}+2\tan^{-1}\left(\tanh\left(\frac{\rho}{2}\right)\right)+\frac{\pi}{2}\right)}.$ (5)
$C_{D}\geq 0$ together with $\lambda\geq 0$ yields the constraints: $0\leq\lambda$ for $\rho\geq 0$; $20\leq\lambda\leq-\frac{\coth(\rho)}{2}\geq 1/$ 2 for $\rho\leq 0$. Note that $\lambda=1/2$ is a critical DGP parameter. Below, we will show $\lambda=1/2$ is a phase-transition point. In the followings, we show only key points. See appendix for detailed derivations.
The vacuum of a strip is dual to AdS soliton in bulk Fujita:2011fp
$ds^{2}=\frac{\frac{dz^{2}}{h(z)}+h(z)d\theta^{2}-dt^{2}+dy^{2}}{z^{2}},$ (6)
where $h(z)=1-z^{3}/z_{h}^{3}$. By using the rescale invariance of AdS, we set $z_{h}=1$. The absence of conical singularities fixes the period of angle $\theta$ in bulk $\beta=4\pi/|h^{\prime}(1)|=4\pi/3$. The strip is defined on the AdS boundary $z=0$ with $0\leq\theta\leq L$, where $L\geq 0$ is the strip width. We derive the energy density $T_{tt}=-1$ from (6) by applying the holographic renormalization deHaro:2000vlm . Compared with the Casimir effect (2), we get the holographic Casimir amplitude
$\kappa_{1}=L^{3},$ (7)
where $L$ can be fixed by NBC (4). For negative brane tension $T\leq 0$, we have
$L_{\text{I}}=\int_{0}^{z_{\max}}\frac{2\Big{(}\sqrt{h\left(z_{\max}\right)}-\lambda h(z)\Big{)}\ dz}{h(z)\sqrt{H(z,z_{\max})-h(z_{\max})}},$ (8)
where $H=\frac{1}{2}h(z)\left(1+2\lambda\sqrt{h\left(z_{\max}\right)}+X\right)$, $X=\sqrt{1+4\lambda^{2}h(z)-4\lambda\sqrt{h\left(z_{\max}\right)}}$ and $z_{\text{max}}$ is the turning point obeying $T=-2\sqrt{h(z_{\text{max}})}$. The case with $T>0$ can be obtained from the complement of the first case $L_{\text{II}}=\beta-L_{\text{I}}\Big{(}T\to-T,\lambda\to-\lambda\Big{)}$. This is the typical trick used in Fujita:2011fp , based on extrinsic curvature flip signs when crossing the brane. In the viewpoint of NBC (4), $T,\lambda$ change signs equivalently. Some comments are in order. First, the integrand of (8) cannot take absolute value, which would miss the case of $T<0$ with $\rho>0$ and cannot produce the correct asymptotical behavior, i.e. $\theta=-\sinh(\rho)z+O(z^{2})$ and $(\theta-L)=\sinh(\rho)z+O(z^{2})$. Second, $L_{\text{I}}$ (8) is negative for sufficiently large $\lambda$. Thus, there is an upper bound of $\lambda$. In fact, the constraint $L\geq 0$ yields $\lambda\leq 1$. Third, the case $20\leq\lambda<1/$ 2 can continuously transform into that of Einstein gravity, while the case $1/2<\lambda\leq 1$ cannot. See Fig. 1 for instance. We name them normal and singular phases, respectively. Fourth, the strip width $L$ decreases with brane tension $T$ Fujita:2011fp and $L\geq 0$ sets an upper bound of $T$ Miyaji:2021ktr . For $T\geq 0$, $L$ should be smaller than angle period $\beta$ to avoid conical singularity. Luckly, $L\leq\beta$ is satisfied automatically for $T\geq 0$ Miyaji:2021ktr . On the other hand, there is no upper bound of $L$ for $T<0$, since the conical singularity is hidden behind the brane Tadashi . Thus, it is irrelevant to the bulk dual of strip. As a result, there is no extra constraint on the lower bound of $T$ and it can take the minimal value $T=-2$ ($\rho\to-\infty$) Miyaji:2021ktr ; Tadashi .
We can derive the ratio $(-\kappa_{1}/C_{D})$ from (5,7,8). See Fig. 1 for example. Fig. 1 (up) denotes the normal phase with $20\leq\lambda<1/$ 2, which shows that the smaller $\rho$, the smaller $(-\kappa_{1}/C_{D})$. Besides, all curves approach the same lower bound (purple curve) from the above in the limit $T\to-2$. On the other hand, Fig. 1 (down) labels the singular phase with $1/2\leq\lambda\leq 1$. It shows that ($-\kappa_{1}/C_{D}$) with $\lambda>1/2$ is always larger than that of Einstein gravity (blue curve). Besides, the minimum of ($-\kappa_{1}/C_{D}$) increases with $\lambda$ and approaches zero when $\lambda\to 1$. Since we are interested in the lower bound of ($-\kappa_{1}/C_{D}$), we focus on the normal phase below.
Figure 1: $(-\kappa_{1}/C_{D})$ in normal phase with $20\leq\lambda<1/$ 2 (up) and singular phase with $1/2<\lambda\leq 1$ (down). The up figure shows all curves approach the holographic limit (purple curve) from the above for $\rho\to-\infty$ ($T\to-2$) in the normal phase. The down figure shows $(-\kappa_{1}/C_{D})$ increases with $\lambda$ for $11/2<\lambda<$ 1. In the limit $\lambda\to 1$, we have $\kappa_{1}/C_{D}\to 0$ in the singular phase. $\lambda=1/2$ is the phase transition point.
Now, we give some analytical discussions. Performing coordinate transformation $z=z_{\text{max}}y$ and expanding (8) around $x=\text{sech}^{2}(\rho)\to 0$, we derive perturbatively for normal phase
$L_{I}=\frac{2\sqrt{\pi}\Gamma\left(\frac{4}{3}\right)}{\Gamma\left(\frac{5}{6}\right)}\Big{(}\frac{(1-2\lambda)^{2}}{x}\Big{)}^{\frac{1}{6}}+O\left(\frac{x}{(1-2\lambda)^{2}}\right)^{\frac{5}{6}}$ (9)
Taking limit $x\to 0$ ($T\to-2$) with $\lambda<1/2$, we obtain a universal limit independent of $\lambda$
$\lim_{T\to-2}(\frac{-\kappa_{1}}{C_{D}})=\lim_{x\to 0}(\frac{-L_{\text{I}}^{3}}{C_{D}})=-\frac{\pi^{5/2}\Gamma\left(\frac{1}{3}\right)^{3}}{108\Gamma\left(\frac{5}{6}\right)^{3}}.$ (10)
It verifies our claim that different holographic models give the same lower bound (3) in the normal phase. Note that (9) implies $L/\beta\to\infty$ for $x\to 0$ ($T\to-2$). Note also that the strip width $L\to Lz_{h}$ can remain finite when recovering $z_{h}\to 0$ in $h(z)$ (6). At the phase transition point $\lambda=1/2$, the lower bound (10) can be achieved by $L=\beta$. See appendix for derivations. The lower bound of $(-\kappa_{1}/C_{D})$ is saturated by negative brane tension. We remark that the negative brane tension is well-defined in AdS/BCFT Fujita:2011fp ; Takayanagi:2011zk ; Miyaji:2021ktr ; Tadashi . First, the brane tension is a cosmological constant rather than a kinetic-energy term on the end-of-the-wolrd (EOW) brane. Of course, the negative cosmological constant is well-defined in AdS/CFT. Second, negative brane tension generally yields a negative A-type boundary central charge. But nothing goes wrong. Recall that the free scalar with the DBC has a negative A-type boundary central charge Jensen:2015swa .
Tests — Let us test the holographic lower bound of the Casimir effect (3) by field theories. First, consider free CFTs. We have $(-\kappa_{1}/C_{D})=-3\pi\zeta(3)/8\approx-1.42$ for free fermions, and $(-\kappa_{1}/C_{D})=-\pi\zeta(3)/2\approx-1.89$ for free scalars with RBC and DBC. Next, consider CFTs with interactions. For Ising model with $\phi^{4}$ interaction at the critical point, we have $2\kappa_{1}\approx 0.820$ and $C_{D}\approx 0.193$ Toldin:2021kun ; Ising referee1 ; Ising referee2 , which yields $(-\kappa_{1}/C_{D})\approx-2.12$. Near the strip boundary, the scalar operator behaves as Cardy:1990xm
$\langle\sigma(z)\rangle=\frac{a_{\sigma}}{(2z)^{\Delta_{\sigma}}}\Big{(}1+B_{\sigma}(\frac{z}{L})^{d}+...\Big{)},$ (11)
where $\Delta_{\sigma}$ is the conformal dimension and $B_{\sigma}$ obeys the universal relation
$\frac{B_{\sigma}}{\Delta_{\sigma}}=2^{d-1}(d-1)\pi^{-\frac{d}{2}}\Gamma\left(\frac{d}{2}\right)\ \frac{\kappa_{1}}{C_{D}}.$ (12)
Here, we use the notation of Toldin:2021kun with $d=3$ for 3d CFTs. For $O(N)$ model with $N=2,3$, Toldin:2021kun obtains $B_{\sigma}\approx 1.21,1.07$ and $\Delta_{\sigma}\approx 0.519088,0.518920$ by applying Monte Carlo simulations. Then (12) gives $(-\kappa_{1}/C_{D})\approx-1.83,-1.62$ for $N=2,3$, respectively. We summarize all the above results in Table.1, which all obey the holographic bound (3). It strongly supports our proposal. Interestingly, (12) suggests the bound of the Casimir effect imposes an upper bound of $B_{\sigma}/\Delta_{\sigma}$. Below we briefly comment on various generalizations of our proposal (3). We show only key points and leave detailed discussions and open questions to future work.
General boundaries — Holography is expected to set a bound of Casimir effect for general boundary shapes, not be limited to strips. Let us provide some evidence. We first consider the wedge space, which is the simplest generalization of a strip. The Casimir effect takes the form
$\langle T^{i}_{\ j}\rangle_{\text{wedge}}=\frac{f(\Omega)}{r^{3}}\text{diag}\Big{(}1,-2,1\Big{)},$ (13)
where $f(\Omega)$ is the Casimir amplitude and $\Omega$ is the opening angle of wedge. In the limit $\Omega\to 0,r\to\infty$ with $r\Omega=L$ fixed, the large $r$ region of the wedge approaches a strip. It results in $f(\Omega)\to\kappa_{1}/\Omega^{3}$ for $\Omega\to 0$ Deutsch:1978sc . On the other hand, we have $f(\Omega)\to(C_{D}\pi/64)(\pi-\Omega)$ in the smooth limit $\Omega\to\pi$ Miao:2024ddp . These two limits suggest
$(\frac{-f(\Omega)}{C_{D}})\geq(\frac{-f(\Omega)}{C_{D}})_{\text{holo}},\ \text{for}\ 0<\Omega\leq\pi,$ (14)
where the holographic bound is derived in the limit $T\to-2$. By applying the method of Miao:2024ddp and this letter, we numerically verify (14) holds. Let us go on to discuss the general boundary. Near any smooth boundary, the normal-normal component of stress tensor takes a universal form in flat space Deutsch:1978sc ; Miao:2017aba
$\langle T_{nn}(z)\rangle=\alpha\frac{\text{Tr}\bar{k}^{2}}{z}-\frac{\alpha}{2}k\text{Tr}\bar{k}^{2}\log(\frac{z}{L})+t_{nn}+O(z)$ (15)
where $z$ is the distance to boundary, $L$ is the system size, and $t_{nn}$ denotes the finite term. The divergent terms appear since we consider the ideal boundary. There is a natural cutoff $z\geq\epsilon$ in real material with $\epsilon$ the lattice length. Remarkably, the displacement operator universally determines the divergent terms, i.e., $\alpha=-C_{D}\pi/16$ Miao:2017aba . It strongly suggests $\langle T_{nn}(\epsilon)\rangle/C_{D}$, or equivalently, $t_{nn}/C_{D}$ has a lower bound for fixed system size and boundary shape. A weaker proposal is for the average pressure
$(\frac{\bar{t}_{nn}}{C_{D}})\geq(\frac{\bar{t}_{nn}}{C_{D}})_{\text{holo}},$ (16)
[…]
Higher dimensions — Let us comment on the holographic bound of Casimir effect for a strip in higher dimensions. The corresponding Casimir amplitude $\kappa_{1}$ and norm of displacement operator $C_{D}$ is given by Miao:2024ddp for Einstein gravity. By taking the limit $T\to-(d-1)$, we derive a universal lower bound
$(\frac{-\kappa_{1}}{C_{D}})\geq\frac{-2^{d-2}d\pi^{d-\frac{1}{2}}\Gamma\left(\frac{d-1}{2}\right)\Gamma\left(\frac{1}{d}\right)^{d}}{\Gamma(d+2)\left(d\Gamma\left(\frac{1}{2}+\frac{1}{d}\right)\right)^{d}}.$ (17)
[…]
Non-CFTs — Interestingly, the lower bound of (3) also works for a general class of non-CFTs. Take an example of a 3d free scalar with mass $m$ and coupling $\xi$ with the Ricci scalar. It is a CFT only if $\xi=1/8$ and $m=0$. For non-CFT parameters, the energy density between parallel plates is no longer a constant Bordag:2009zz . What is worse, it is divergent near the boundary. Fortunately, the total energy per area is finite Bordag:2009zz . On the other hand, the pressure is still a finite constant. Thus, we define $\kappa_{1}=\langle T_{nn}(0)\rangle L^{3}/(-2)$ generally. By applying the textbook method, we obtain for both RBC and DBC, i.e., $16\pi\kappa_{1}=-2M^{2}\log\left(1-e^{-2M}\right)+2M\text{Li}_{2}\left(e^{-2M}\right)+\text{Li}_{3}\left(e^{-2M}\right)$, where $M=mL>0$ is a dimensionless parameter, and Li denotes the Polylogarithmic function. Remarkably, $\kappa_{1}$ is independent of $\xi$ and decreases with mass
$\frac{d\kappa_{1}}{dM}=\frac{M^{2}(1-\coth(M))}{8\pi}<0.$ (18)
It means the Casimir effect is suppressed by mass Bordag:2009zz . For non-CFTs, we can define $C_{D}$ via $\langle D(y)D(0)\rangle=C_{D}/|y|^{6}+O(1/|y|^{5})$ in half space, where $D=T_{nn}(0)$ Billo:2016cpy . At small distance $m|y|\ll 1$, the mass effect can be ignored. As a result, $C_{D}$ is independent of mass. The above discussions show $(-\kappa_{1}/C_{D})$ is larger than that of massless scalar, thus obeys the holographic lower bound (3). The above arguments apply to general massive free field theories.
Discussions — In conclusion, we propose that holography sets a fundamental bound for the Casimir amplitude ratio to the norm of displacement operator. The Casimir amplitude is generally defined as the average pressure $\bar{t}_{nn}$ on the boundary. We take 3d CFTs in a strip as an example and then briefly discuss the generalization to other boundary shapes, higher dimensions, and non-CFTs. It is interesting to test our proposal in experiments of critical systems Krech:Casimir Effect . Finding a field-theoretical proof or counterexample for the holographic bound of the Casimir effect is also enjoyable. Usually, a bound is set by either free or strongly coupled theory. For entanglement entropy (EE), there is a similar bound set by free theory, i.e., $(-F_{0}/C_{T})\geq(-F_{0}/C_{T})_{\text{free scalar}}$ Bueno:2023gey , where $F_{0}$ is the universal coefficient of EE and $C_{T}$ is a bulk central charge. We remark that the dual brane for EE is tensionless while the one for the Casimir effect is tensive. With a tensive brane, we have more parameters to get a smaller lower bound of Casimir effect in holography than in free theory. Ising model of Table. 1 also shows a free scalar does not set the lower bound of Casimir effect. In AdS3, conformal bootstrap restricts brane tension $|T|\leq 0.99$ sightly stronger than $|T|\leq 1$ Collier:2021ngi . If this is also the case for both $T$ and DGP coupling $\lambda$ in higher dimensions, it may give a slightly stronger lower bound of the Casimir effect. Anyway, this letter takes a significant step toward exploring the lower bound of the Casimir effect. Many open problems remain to be explored.
Acknowledgements
We thank T. Takayanagi, J. X. Lu for valuable comments and discussions. We are grateful to the conference “Gauge Gravity Duality 2024”, where the work was completed. This work is supported by the National Natural Science Foundation of China (No.12275366).
Appendix A Bound of Casimir effect from DGP gravity
This appendix studies the holographic bound of Casimir effect in DGP gravity. In particular, we give detailed derivations of eqs.(8, 10) in the main text.
Let us quickly recall some key points. The vacuum of a strip is dual to AdS soliton
$ds^{2}=\frac{\frac{dz^{2}}{h(z)}+h(z)d\theta^{2}-dt^{2}+dy^{2}}{z^{2}},$ (19)
where $h(z)=1-z^{3}/z_{h}^{3}$, and $\theta$ has the range $0\leq\theta\leq L$ on the AdS boundary $z=0$. We label the embedding function of brane $Q$ as
$\theta=S(z).$ (20)
It obeies NBC
$K^{ij}-(K-T+\lambda\mathcal{R})h^{ij}+2\lambda\mathcal{R}^{ij}=0,$ (21)
where the tension is parameterized as Miao:2023mui
$T=2\tanh(\rho)-2\lambda\text{sech}^{2}(\rho).$ (22)
This parameterization can be derived from the gravity dual of half space. Since the region near any smooth boundary can be approximately by half space at leading order of $z$, it works for general boundaries. Substituting (20) into NBC (21), we derive near the two boundaries of strip
$\theta=-\sinh(\rho)z+O(z^{2}),\ \ \theta-L=\sinh(\rho)z+O(z^{2}).$ (23)
Let us first study the case with negative brane tension $T\leq 0$, the other case with $T>0$ can be obtained by the complement of the first case. This is the typical trick used in Fujita:2011fp . Let us explain more. See Fig. 2 for the geometry of holographic dual of strip. The region between red/blue curves (branes) and black line is the bulk dual of strip I with $T\leq 0$; its complement in bulk is the gravity dual for strip II (green line) with $T>0$. As shown in Fig. 2, the gravity duals of strip I and strip II shares the same EOW branes (blue or red curve depending on the theory parameter). As a result, the NBCs (21) for strip I and strip II take the same values. However, the extrinsic curvatures $K_{ij}$ flip signs, while the induced metric $h_{ij}$ and intrinsic Ricci tensor $\mathcal{R}_{ij}$ remain invariant when crossing the branes. In the viewpoint of NBC (21), $T$ and $\lambda$ equivalently change signs when crossing the branes. Note that the bulk dual for strip II contains the ‘horizon’ $z=z_{h}$. To remove the conical singularity at $z=z_{h}$, we fix the period of angle $\theta$ as $\beta=4\pi/|h^{\prime}(z_{h})|=4\pi z_{h}/3$. Thus for $T>0$, the left and right vertical dotted lines of Fig. 2 are identified due to the periodicity of angle $\theta$. On the other hand, for strip I with $T\leq 0$, the ‘horizon’ and potential conical singularity is hidden behind the branes (red/blue curves). Thus, the conical singularity is irrelevant and $\theta$ can be non-periodic for $T\leq 0$.
Without loss of generality, we focus on the strip I with $T\leq 0$ and $z_{h}=1$ below. According to (23), the blue and red curves of Fig. 2 correspond to $\rho<0$ and $\rho>0$, respectively. We focus on the left half of branes, the right half can be obtained by symmetry. NBC (21) yields an independent equation for $T\leq 0$
$T=-\frac{2h(z)\left(h(z)S^{\prime}(z)\sqrt{h(z)S^{\prime}(z)^{2}+\frac{1}{h(z)}}+\lambda\right)}{h(z)^{2}S^{\prime}(z)^{2}+1}.$ (24)
Let us verify that it corresponds to the left half of branes. From (22,24), we derive $S^{\prime}(0)=-\sinh(\rho)$ near $\theta_{1}$ of Fig. 2, which agrees with the first equation of (23). As shown in Fig. 2, $S^{\prime}(z_{\text{max}})=\infty$ and $S^{\prime}(z_{c})=0$ at the turning points $z_{\text{max}}$ and $z_{c}$ for the left half of branes. Substituting $S^{\prime}(z_{\text{max}})=\infty$ and $S^{\prime}(z_{c})=0$ into (24), we obtain
$T=-2\sqrt{h(z_{\text{max}})}=-2\lambda h(z_{c}).$ (25)
Note that the turning point $z_{c}$ appears only in the red curve with $\rho>0$. Solving (24) and (25), we get
$S^{\prime}(z)=\pm\frac{\sqrt{h\left(z_{\max}\right)}-\lambda h(z)}{h(z)\sqrt{H(z,z_{\max})-h(z_{\max})}},$ (26)
where $H=\frac{1}{2}h(z)\left(1+2\lambda\sqrt{h\left(z_{\max}\right)}+X\right)$, $X=\sqrt{1+4\lambda^{2}h(z)-4\lambda\sqrt{h\left(z_{\max}\right)}}$. We choose the positive sign, i.e., $S^{\prime}(z)\sim\Big{(}\sqrt{h\left(z_{\max}\right)}-\lambda h(z)\Big{)}$, for the left halves of red and blue curves to get the correct behavior $S^{\prime}(0)=-\sinh(\rho)$ near $\theta_{1}$ of Fig. 2. The negative sign of (26) corresponds to the right halves of branes. From (20,26), we derive the width of strip I for $T\leq 0$
$L_{\text{I}}=\int_{0}^{z_{\max}}\frac{2\Big{(}\sqrt{h\left(z_{\max}\right)}-\lambda h(z)\Big{)}\ dz}{h(z)\sqrt{H(z,z_{\max})-h(z_{\max})}},$ (27)
where $z_{\max}=(1-T^{2}/4)^{1/3}$ from (25). We stress that the above formula works for both the cases of blue curve and red curve of Fig. 2. In particular, the integrand $2S^{\prime}(z)\sim 2\Big{(}\sqrt{h\left(z_{\max}\right)}-\lambda h(z)\Big{)}$ flips signs at turning point $z_{c}$ (25), which is the expected feature for the red curve. It suggests the integrand cannnot be chosen as the absolute value $2|S^{\prime}(z)|\sim 2\Big{|}\sqrt{h\left(z_{\max}\right)}-\lambda h(z)\Big{|}$, which would miss the case of red curve.
Figure 2: Geometry of holographic strip: a portion of AdS soliton. The region between red/blue curves (branes) and black line is the bulk dual of strip I with negative brane tension $T\leq 0$; its complement in bulk is the gravity dual for strip II (green line) with $T>0$. The gravity dual of strip II contains the ‘horizon’ $z=z_{h}$. To remove the conical singularity on it, the angle $\theta$ should be periodic. Thus green lines for strip II are connected. Without loss of generality, we focus on strip I with $T\leq 0$. We have $(\rho<0)$ and $(\rho>0)$ for the blue and red curves, respectively. $z_{\text{max}}$ and $z_{c}$ denote the turning points with $\theta^{\prime}(z_{\text{max}})=\infty$ and $\theta^{\prime}(z_{c})=0$.
Recall that strip II is the complement of strip I and $(T,\lambda)$ of NBC (21) flip signs when crossing the brane, we get the width of strip II for $T\geq 0$
$L_{\text{II}}=\beta-L_{\text{I}}\Big{(}T\to-T,\lambda\to-\lambda\Big{)},$ (28)
where, as we discuss above, the angle period $\beta=4\pi z_{h}/3$ is meaningful only for $T\geq 0$. In total, we have
$L=\begin{cases}\ L_{\text{I}},&\ \text{for}\ T\leq 0,\\ \ L_{\text{II}},&\ \text{for}\ T\geq 0.\end{cases}$ (29)
One can check $L$ is continuous at $T=0$, which is a test of our results. See Fig. 3 for example. As mentioned in main text, the constraint $0\leq L$ set an upper bound of $\lambda$. It is clear from (27) that large $\lambda$ would give negative $L$. We numerically get the range $0\leq\lambda\leq 1$ for $L\geq 0$. As shown in Fig. 3 , the case $20\leq\lambda\leq 1/$ 2 can continuously transform into that of Einstein gravity, while the case $1/2<\lambda\leq 1$ cannot. We name them normal and singular phases, respectively. $\lambda=1/2$ is the phase-transition point.
Figure 3: Strip widths in normal phase with $20\leq\lambda<1/$ 2 (up) and in singular phase with $1/2<\lambda\leq 1$ (down). Note that $L$ can be larger than $\beta$ for negative enough $\rho$ in the normal phase. We show only range of $L\leq\beta$ for simplicity. On the other hand, $L$ is always smaller than $\beta$ in the singular phase. In both cases, $0\leq L$ imposes an upper bound of $\rho$.
Now we make some analytical discussions. Performing coordinate transformation $z=z_{\text{max}}y$ and expanding (27) around $x=\text{sech}^{2}(\rho)\to 0$, we derive perturbatively for normal phase
[…]
$=$ $\frac{2\sqrt{\pi}\Gamma\left(\frac{4}{3}\right)}{\Gamma\left(\frac{5}{6}\right)}\Big{(}\frac{(1-2\lambda)^{2}}{x}\Big{)}^{\frac{1}{6}}+(1-2\lambda)\ O\left(\frac{x}{(1-2\lambda)^{2}}\right)^{\frac{5}{6}}$
The norm of displacement operator becomes for $\rho<0$
$C_{D}=\frac{32}{\frac{2\pi\lambda\sqrt{x}}{1-2\lambda\sqrt{1-x}}+\pi\cot^{-1}\left(\sqrt{\frac{1}{x}-1}\right)}.$ (31)
Taking the limit $x\to 0$ ($\rho\to-\infty$) with $\lambda<1/2$, we obtain a universal limit
$\lim_{\rho\to-\infty}(\frac{-\kappa_{1}}{C_{D}})=\lim_{x\to 0}(\frac{-L_{\text{I}}^{3}}{C_{D}})=-\frac{\pi^{5/2}\Gamma\left(\frac{1}{3}\right)^{3}}{108\Gamma\left(\frac{5}{6}\right)^{3}},$ (32)
in normal phase. Note that $L_{\text{I}}$ (A) goes to infinity for $x\to 0$. Recovering $z_{h}$ in $h(z)=1-z^{3}/z_{h}^{3}$, we can keep the strip width $L\to z_{h}L$ finite in the limit $x\to 0$ and $z_{h}\to 0$. Let us go on to discuss the case at the phase transition point $\lambda\to 1/2$. For our purpose, we consider the limit $x\to 0$, $\lambda\to 1/2$ with $x/(1-2\lambda)^{2}$ finite. In such limit, (A) becomes
$L_{\text{I}}\to\frac{2\sqrt{\pi}\Gamma\left(\frac{4}{3}\right)}{\Gamma\left(\frac{5}{6}\right)}\Big{(}\frac{(1-2\lambda)^{2}}{x}\Big{)}^{\frac{1}{6}}.$ (33)
Interestingly, the second term of (A) vanishes for $\lambda\to 1/2$. In fact, this is also the case for higher-order expansions. From (33) and $L_{\text{I}}=\beta=4\pi/3$, we solve
$x=\frac{729\Gamma\left(\frac{4}{3}\right)^{6}}{64\pi^{3}\Gamma\left(\frac{5}{6}\right)^{6}}(1-2\lambda)^{2}+O(1-2\lambda)^{3}.$ (34)
Substituting (34) into (31), we get
$C_{D}=\frac{256\sqrt{\pi}\Gamma\left(\frac{5}{6}\right)^{3}}{27\Gamma\left(\frac{4}{3}\right)^{3}}+O(\lambda-\frac{1}{2}).$ (35)
Finally, we obtain
$\lim_{\lambda\to 1/2,L=\beta}(\frac{-\kappa_{1}}{C_{D}})=\lim_{\lambda\to 1/2}(\frac{-\beta^{3}}{C_{D}})=-\frac{\pi^{5/2}\Gamma\left(\frac{1}{3}\right)^{3}}{108\Gamma\left(\frac{5}{6}\right)^{3}},$ (36)
which reproduces the universal limit (32). Interestingly, $L=\beta$ instead of $L/\beta\to\infty$ at the phase transition point $\lambda\to 1/2$.

### v3

@@
 Abstract
-Inspired by the Kovtun-Son-Starinet bound, we propose that holography impose a lower bound on the Casimir effect.
-For simplicity, we focus on the Casimir effect between parallel planes for 3d conformal field theories and briefly comment on the generalizations to other boundary shapes and higher dimensions.
-Remarkably, the ghost-free holographic models impose a universal lower bound of the Casimir effect.
-We verify the holographic bound by free theories, Ising model, and $O(N)$ model with $N=2,3$ at critical points.
-Remarkably, the holographic bound is also obeyed by a general class of quantum field theories without conformal symmetries.
-It is interesting to find a field-theoretical proof or counterexample for the holographic bound of Casimir effect.
-Introduction — Casimir effect Casimir:1948dh is a novel quantum effect that originates from the changes of vacuum fluctuations due to the boundary Plunien:1986ca ; Bordag:2001qi ; Milton:2004ya ; Bordag:2009zz .
-It is a candidate of dark energy Wang:2016och and plays a vital role in QCD Vepstas:1984sw and worm-hole physics Morris:1988tu ; Maldacena:2020sxe .
+Inspired by the Kovtun-Son-Starinet bound, we propose that holography imposes a lower bound on the Casimir effect.
+For simplicity, we focus on the Casimir effect between parallel planes for three-dimensional conformal field theories and briefly comment on the generalizations to other boundary shapes and higher dimensions.
+Remarkably, the ghost-free holographic models impose a universal lower bound on the Casimir effect.
+We verify the holographic bound by free theories, the Ising model, and $O(N)$ models with $N=2,3$ at critical points and prove it for the two-dimensional case.
+Remarkably, a general class of quantum field theories without conformal symmetries also obeys the holographic bound.
+Introduction — Casimir effect Casimir:1948dh is an interesting quantum effect that originates from the changes of vacuum fluctuations due to the boundary Plunien:1986ca ; Bordag:2001qi ; Milton:2004ya ; Bordag:2009zz .
+It is a candidate of dark energy Wang:2016och and plays a vital role in QCD Vepstas:1984sw and wormhole physics Morris:1988tu ; Maldacena:2020sxe .
 It has been measured in experiments Mohideen:1998iz ; Bressi:2002fr ; Klimchitskaya:2009cw and has significant potential applications in nanotechnology.
 As the vacuum energy, Casimir energy is defined as the lowest energy of a quantum system with a boundary.
-Since energy should be bounded from the below, it is expected that there is a fundamental lower bound of the Casimir effect.
+Since energy should be bounded from below, it is expected that there is a fundamental lower bound of the Casimir effect.
 This letter explores this vital problem.
@@
 where $\hbar$ is the reduced Planck constant and $k_{B}$ is the Boltzmann constant.
-The lower bound is obtained from the strongly-coupled CFT dual to Einstein gravity Policastro:2001yc and can be modified slightly by higher derivative gravity Brigante:2007nu ; Kats:2007mq ; Brigante:2008gz .
-Now, let us explore whether AdS/CFT set a similar lower bound for the Casimir effect.
-For simplicity, we focus on the Casimir effect for 3d CFTs between parallel planes (strip) and comment on various generalizations at the end of the letter.
+The lower bound is obtained from the strongly coupled CFT dual to Einstein gravity Policastro:2001yc and can be modified slightly by higher derivative gravity Brigante:2007nu ; Kats:2007mq ; Brigante:2008gz .
+Now, let us explore whether AdS/CFT sets a similar lower bound for the Casimir effect.
+For simplicity, we focus on the Casimir effect for three-dimensional (3D) CFTs between parallel planes (strip) and comment on various generalizations at the end of the Letter.
 We consider flat space and impose the same boundary conditions on the two planes.
-Then, the Casimir effect takes a universal form
+Then, the Casimir effect takes a universal form Miao:2024ddp
 $\langle T^{i}_{\ j}\rangle_{\text{strip}}=\kappa_{1}\frac{\hbar c}{L^{3}}\text{diag}\Big(1,-2,1\Big),$ (2)
-where $\kappa_{1}$ is the dimensionless Casimir amplitude, $c$ is the speed of light, $L$ is the distance between the two plates.
-For a CFT with fixed strip width $L$, $(-\kappa_{1})$ should be bounded from below to avoid negative infinity energy density $(T_{tt}=-\kappa_{1}\hbar c/L^{3})$ and pressure $(T_{nn}=-2\kappa_{1}\hbar c/L^{3})$.
-Note that the Casimir amplitude $\kappa_{1}$ is an extensive parameter in the sense that if one CFT has a Casimir amplitude $\kappa_{1}$, then $N$-copies of the same CFT has a Casimir amplitude $N\kappa_{1}$.
-It implies only $(-\kappa_{1})$ over another extensive parameter could have a lower bound.
-Natural candidates of these extensive parameters are the central charges of CFTs.
-Since the Casimir effect is a boundary effect, the most natural candidates are boundary central charges defined in Weyl anomalyJensen:2015swa $\mathcal{A}=\int_{\partial M}d^{2}x\sqrt{|\sigma|}\Big{(}aR_{\partial M}+b\text{tr}\bar{k}^{2}\Big{)},$ where $R_{\partial M}$ and $\bar{k}$ are the intrinsic Ricci scalar and traceless parts of extrinsic curvatures on the boundary $\partial M$, respectively.
-$a$ is the A-type boundary central charge, which decreases under the boundary RG flow Jensen:2015swa .
+where $\kappa_{1}$ is the dimensionless Casimir amplitude, $c$ is the speed of light, and $L$ is the strip width.
+For the fixed $L$, $\kappa_{1}$ should be bounded from below to avoid negative infinity energy density $T_{tt}=-\kappa_{1}\hbar c/L^{3}$ and pressure $T_{nn}=-2\kappa_{1}\hbar c/L^{3}$.
+Note that the Casimir amplitude $\kappa_{1}$ is an extensive parameter in the sense that if one CFT has a Casimir amplitude $\kappa_{1}$, then $N$-copies of the same CFT have a Casimir amplitude $N\kappa_{1}$.
+It implies that only $\kappa_{1}$ over another extensive parameter could have a lower bound.
+Natural candidates of these extensive parameters are boundary central charges in Weyl anomalyJensen:2015swa $\mathcal{A}=\int_{\partial M}d^{2}x\sqrt{|\sigma|}\Big(aR_{\partial M}+b\text{tr}\bar{k}^{2}\Big)$, where $R_{\partial M}$ and $\bar{k}$ are the intrinsic Ricci scalar and traceless parts of extrinsic curvatures on the boundary $\partial M$, respectively.
+$a$ is the A-type boundary central charge, which decreases under the boundary renomalization group (RG) flow Jensen:2015swa .
 We remark that $a$ can be positive, zero, or negative.
-For example, $a=\pm 1/(384\pi)$ for conformally coupled free scalar with Robin boundary condition (RBC) and Dirichlet boundary condition (DBC) Jensen:2015swa , respectively; $a\sim\sinh(\rho)$ with $-\infty<\rho<\infty$ for holographic CFT3 Fujita:2011fp .
-Since $a$ can be zero and negative, there is no lower bound for $(-\kappa_{1}/a)$.
-On the other hand, the B-type boundary central charge $b$ is always positive since it is related to the Zamolodchikov norm of displacement operator, i.e., $b=C_{D}\pi/32>0$.
+For example, $a=\pm 1/(384\pi)$ for conformally coupled free scalar with Robin boundary condition (RBC) and Dirichlet boundary condition (DBC) Jensen:2015swa , respectively; $a\sim\sinh(\rho)$ with $\infty<\rho<\infty$ for holographic CFT3 Fujita:2011fp .
+Since $a$ can be zero and negative, there is no lower bound for $\kappa_{1}/a$.
+On the other hand, the B-type boundary central charge $b$ is always positive since it is related to the Zamolodchikov norm of the displacement operator, i.e., $b=C_{D}\pi/32>0$.
 The displacement operator $D(y)$ describes the violation of space translation invariance normal to the boundary Billo:2016cpy , i.e., $\nabla_{i}T^{ij}=\delta(x)D(y)n^{j}$.
-The two-point function defines its Zamolodchikov norm, i.e., $\langle D(y)D(0)\rangle=C_{D}/|y|^{6}$.
-By definition, $C_{D}=32b/\pi$ is always positive.
-Thus, we propose the ratio of Casimir amplitude to norm of displacement operator has a lower bound set by holography
+The two-point function defines its Zamolodchikov norm, i.e., $\langle D(y)D(0)\rangle=C_{D}c^{2}\hbar^{2}/|y|^{6}$.
+By definition, $C_{D}$ is dimensionless and positive.
+The above arguments suggest $\kappa_{1}/C_{D}$ should have a physical lower bound.
+Usually, a bound is set either by free theories or strongly coupled CFTs dual to gravity.
+We have checked all the known examples and find that, similar to the KSS bound, holography gives the smallest $\kappa_{1}/C_{D}$.
+Thus, we propose that the ratio of Casimir amplitude to the norm of the displacement operator has a lower bound set by holography
+$-\frac{\kappa_{1}}{C_{D}}\geq-\frac{\pi^{5/2}\Gamma\left(\frac{1}{3}\right)^{3}}{108\Gamma\left(\frac{5}{6}\right)^{3}}\approx-2.17,\ \ \ \text{for 3D CFTs}.$ (3)
+It means the energy density $\kappa_{1}/L^{3}$ per degree of freedom $C_{D}$ is bounded from below for a physical system with fixed size $L$.
+Remarkably, unlike the KSS bound (1), different holographic models give the same universal lower bound (3) for the Casimir effect.
+We verify the lower bound (3) by free CFTs, the Ising model, and the $O(N)$ model with $N=2,3$ at critical points.
+Furthermore, we prove the corresponding bound for two-dimensional (2D) CFTs, which strongly supports our general proposal.
+Below, we discuss the test of our proposal, the derivation of the holographic bound, and its various generalizations.
+For simplicity, we focus on natural units with $c=\hbar=k_{B}=1$ in the following.
+Table 1: $(-\kappa_{1}/C_{D})$ for various 3D CFTs
+Dirac fermion Scalar Ising O(2) O(3) AdS/BCFT
+1.42-1.42 1.89-1.89 -2.12 -1.83 -1.62 2.17-2.17
+Tests — Let us test the holographic bound (3) by various 3D CFTs, which is summarized in Table 1.
+We have $\kappa_{1}/C_{D}=-3\pi\zeta(3)/8\approx-1.42$ for Dirac fermions with bag boundary condition (BC) Bellucci:2009hh ; McAvity:1993ue , and $\kappa_{1}/C_{D}=-\pi\zeta(3)/2\approx-1.89$ for free scalars with RBC and DBC Romeo:2000wt ; Miao:2018dvm .
+For Ising model with $\phi^{4}$ interaction at the critical point, we have $2\kappa_{1}\approx 0.820$ and $C_{D}\approx 0.193$ Toldin:2021kun ; Ising referee1 ; Ising referee2 , which yields $\kappa_{1}/C_{D}\approx-2.12$.
+Here, we impose with $(++)$ BC for the Ising model, where the spins on the two boundaries have the same signs.
+Near the strip boundary, the scalar operator behaves as Cardy:1990xm
 […]
-Remarkably, unlike the KSS bound (1), different holographic models give the same universal lower bound (3) for Casimir effect.
-We verify the lower bound (3) by free CFTs, Ising model, and $O(N)$ model with $N=2,3$ at critical points.
-See Table.
-1.
-Below we discuss the derivation of the holographic lower bound and its various tests and generalizations.
-We focus on natural units with $c=\hbar=k_{B}=1$ in the followings.
-Table 1: $(-\kappa_{1}/C_{D})$ for various 3d CFTs
-Fermion Scalar Ising O(2) O(3) Holography
-$-1.42$ $-1.89$2.12 -1.83 -1.62 $-2.17$
+where $\Delta_{\sigma}$ is the conformal dimension and $B_{\sigma}$ obeys the universal relation
+$\frac{B_{\sigma}}{\Delta_{\sigma}}=2^{d-1}(d-1)\pi^{-\frac{d}{2}}\Gamma\left(\frac{d}{2}\right)\ \frac{\kappa_{1}}{C_{D}}.$ (5)
+Here, we use the notation of Toldin:2021kun with $d=3$ for 3D CFTs.
+For $O(N)$ model with $N=2,3$ and open BCs, Toldin:2021kun obtains $B_{\sigma}\approx 1.21,1.07$ and $\Delta_{\sigma}\approx 0.519088,0.518920$ by applying Monte Carlo simulations.
+Then (5) gives $\kappa_{1}/C_{D}\approx-1.83,-1.62$ for $N=2,3$, respectively.
+We summarize all the above results in the Table 1, which all obey the holographic bound (3) will be derived below.
+It strongly supports our proposal.
+Interestingly, (5) suggests the bound of the Casimir effect imposes an upper bound of $B_{\sigma}/\Delta_{\sigma}$.
 Holographic bound — Let us study the holographic bound of the Casimir effect of a strip in AdS/BCFT Takayanagi:2011zk .
-We are interested in the general ghost-free gravity in AdS4, which turns out to be Dvali-Gabadadze-Porrati (DGP) gravity Dvali:2000hr $I=\int_{N}d^{4}x\sqrt{|g|}\Big{(}R+6\Big{)}+2\int_{Q}d^{3}y\sqrt{|h|}\Big{(}K-T+\lambda\mathcal{R}\Big{)}$, where $K$ is the extrinsic curvature, $T$ is the brane tension, $\lambda$ is the DGP parameter, $R$ and $\mathcal{R}$ are Ricci scalars in bulk $N$ and on the brane $Q$ respectively.
-We have set Newton’s constant $116\pi G_{N}=$ 1 and AdS radius $l=1$ for simplicity.
+We are interested in the general ghost-free gravity in AdS4, which turns out to be Dvali-Gabadadze-Porrati (DGP) gravity,Dvali:2000hr
+$I=\int_{N}d^{4}x\sqrt{|g|}\Big(R+6\Big)+2\int_{Q}d^{3}y\sqrt{|h|}\Big(K-T+\lambda\mathcal{R}\Big),$ (6)
+where $K$ is the extrinsic curvature, $T$ is the brane tension, $\lambda$ is the DGP parameter, $R$ and $\mathcal{R}$ are Ricci scalars in bulk $N$ and on the brane $Q$, respectively.
+We have set Newton’s constant $16\pi G_{N}=1$ and AdS radius $l=1$ for simplicity.
 Note that higher derivative gravity and negative DGP gravity generally suffer the ghost problems Miao:2023mui , while Gauss-Bonnet/Lovelock gravity appears only in dimensions higher than four.
 Thus, positive DGP gravity with $\lambda\geq 0$ is the general ghost-free gravity theory in AdS4.
-Following Takayanagi:2011zk , we impose Neumann boundary condition (NBC) on the brane $Q$
+The case $\lambda=0$ corresponds to Einstein gravity.
+Following Takayanagi:2011zk , we impose the Neumann boundary condition (NBC) on the brane $Q$
 $K^{ij}-(K-T+\lambda\mathcal{R})h^{ij}+2\lambda\mathcal{R}^{ij}=0,$ (7)
-where $T=2\tanh(\rho)-2\lambda\text{sech}^{2}(\rho)$ is a useful parameterization Miao:2023mui . The holographic norm of the displacement operator is given by Miao:2023mui ; Miao:2018dvm
+where $T=2\tanh(\rho)-2\lambda\text{sech}^{2}(\rho)$ is a useful parametrization Miao:2023mui . The holographic norm of the displacement operator is given by Miao:2023mui ; Miao:2018dvm
 $C_{D}=\frac{32}{\pi\left(\frac{2\lambda}{2\lambda\sinh(\rho)+\cosh(\rho)}+2\tan^{-1}\left(\tanh\left(\frac{\rho}{2}\right)\right)+\frac{\pi}{2}\right)}.$ (8)
-$C_{D}\geq 0$ together with $\lambda\geq 0$ yields the constraints: $0\leq\lambda$ for $\rho\geq 0$; $20\leq\lambda\leq-\frac{\coth(\rho)}{2}\geq 1/$ 2 for $\rho\leq 0$.
-Note that $\lambda=1/2$ is a critical DGP parameter.
-Below, we will show $\lambda=1/2$ is a phase-transition point.
-In the followings, we show only key points.
-See appendix for detailed derivations.
-The vacuum of a strip is dual to AdS soliton in bulk Fujita:2011fp
+$C_{D}\geq 0$ together with $\lambda\geq 0$ yields the constraints: $0\leq\lambda$ for $\rho\geq 0$; $0\leq\lambda\leq-\frac{\coth(\rho)}{2}\geq 1/2$ for $\rho\leq 0$.
+Below, we will show that $\lambda=1/2$ is a phase-transition point and there are other physical constraints for $\lambda$.
+In the following, we show only key points and leave the detailed derivations to the supplemental material.
+The vacuum of a strip is dual to the AdS soliton Fujita:2011fp
 $ds^{2}=\frac{\frac{dz^{2}}{h(z)}+h(z)d\theta^{2}-dt^{2}+dy^{2}}{z^{2}},$ (9)
 where $h(z)=1-z^{3}/z_{h}^{3}$.
-By using the rescale invariance of AdS, we set $z_{h}=1$.
-The absence of conical singularities fixes the period of angle $\theta$ in bulk $\beta=4\pi/|h^{\prime}(1)|=4\pi/3$.
+The absence of conical singularities fixes the period of angle $\theta$ in bulk $\beta=4\pi/|h^{\prime}(z_{h})|=4z_{h}\pi/3$.
 The strip is defined on the AdS boundary $z=0$ with $0\leq\theta\leq L$, where $L\geq 0$ is the strip width.
-We derive the energy density $T_{tt}=-1$ from (6) by applying the holographic renormalization deHaro:2000vlm .
+We derive the energy density $T_{tt}=-1/z_{h}^{3}$ from (9) by applying the holographic renormalization deHaro:2000vlm .
 Compared with the Casimir effect (2), we get the holographic Casimir amplitude
-$\kappa_{1}=L^{3},$ (7)
-where $L$ can be fixed by NBC (4). For negative brane tension $T\leq 0$, we have
+$\kappa_{1}=L^{3}/z_{h}^{3},$ (10)
+where $L$ can be fixed by NBC (7). By using the rescale invariance of AdS, we set $z_{h}=1$ below. For negative brane tension $T\leq 0$, we derive the strip width
 $L_{\text{I}}=\int_{0}^{z_{\max}}\frac{2\Big(\sqrt{h\left(z_{\max}\right)}-\lambda h(z)\Big)\ dz}{h(z)\sqrt{H(z,z_{\max})-h(z_{\max})}},$ (11)
 where $H=\frac{1}{2}h(z)\left(1+2\lambda\sqrt{h\left(z_{\max}\right)}+X\right)$, $X=\sqrt{1+4\lambda^{2}h(z)-4\lambda\sqrt{h\left(z_{\max}\right)}}$ and $z_{\text{max}}$ is the turning point obeying $T=-2\sqrt{h(z_{\text{max}})}$.
-The case with $T>0$ can be obtained from the complement of the first case $L_{\text{II}}=\beta-L_{\text{I}}\Big{(}T\to-T,\lambda\to-\lambda\Big{)}$.
-This is the typical trick used in Fujita:2011fp , based on extrinsic curvature flip signs when crossing the brane.
+The strip width with $T>0$ can be obtained from the complement of that with $T<0$ by $L_{\text{II}}=\beta-L_{\text{I}}\Big(T\to-T,\lambda\to-\lambda\Big)$.
+It is the typical trick used in Fujita:2011fp , based on extrinsic curvature flip signs when crossing the brane.
 In the viewpoint of NBC (7), $T,\lambda$ change signs equivalently.
 Some comments are in order.
-First, the integrand of (8) cannot take absolute value, which would miss the case of $T<0$ with $\rho>0$ and cannot produce the correct asymptotical behavior, i.e.
-$\theta=-\sinh(\rho)z+O(z^{2})$ and $(\theta-L)=\sinh(\rho)z+O(z^{2})$.
-Second, $L_{\text{I}}$ (8) is negative for sufficiently large $\lambda$.
-Thus, there is an upper bound of $\lambda$.
-In fact, the constraint $L\geq 0$ yields $\lambda\leq 1$.
-Third, the case $20\leq\lambda<1/$ 2 can continuously transform into that of Einstein gravity, while the case $1/2<\lambda\leq 1$ cannot.
+First, $L_{\text{I}}$ (11) is negative for sufficiently large $\lambda$.
+The physical constraint $L\geq 0$ yields $\lambda\leq 1$.
+Second, the case $0\leq\lambda<1/2$ can continuously transform into that of Einstein gravity, while the case $1/2<\lambda\leq 1$ cannot.
 See Fig.
@@
 We name them normal and singular phases, respectively.
-Fourth, the strip width $L$ decreases with brane tension $T$ Fujita:2011fp and $L\geq 0$ sets an upper bound of $T$ Miyaji:2021ktr .
-For $T\geq 0$, $L$ should be smaller than angle period $\beta$ to avoid conical singularity.
-Luckly, $L\leq\beta$ is satisfied automatically for $T\geq 0$ Miyaji:2021ktr .
+Third, the strip width $L$ decreases with brane tension $T$ Fujita:2011fp and $L\geq 0$ sets an upper bound of $T$ Miyaji:2021ktr .
+For $T\geq 0$, $L$ should be smaller than the angle period $\beta$ to avoid conical singularity.
+Luckily, $L\leq\beta$ is satisfied automatically for $T\geq 0$ Miyaji:2021ktr .
 On the other hand, there is no upper bound of $L$ for $T<0$, since the conical singularity is hidden behind the brane Tadashi .
-Thus, it is irrelevant to the bulk dual of strip.
-As a result, there is no extra constraint on the lower bound of $T$ and it can take the minimal value $T=-2$ ($\rho\to-\infty$) Miyaji:2021ktr ; Tadashi .
-We can derive the ratio $(-\kappa_{1}/C_{D})$ from (5,7,8).
+Thus, it is irrelevant to the bulk dual of the strip.
+As a result, there is no extra constraint on the lower bound of $T$, and it can take the minimal value T=2T=-2 Miyaji:2021ktr ; Tadashi .
+We remark that $T\geq-2$ sets no constraint on $\rho$ in the normal phase, while it sets a lower bound of $\rho$ in the singular phase.
 See Fig.
-1 for example.
+1 (bottom).
+We can derive the ratio $\kappa_{1}/C_{D}$ from (8,10,11).
+Refer to Fig.
+1 for details.
 Fig.
-1 (up) denotes the normal phase with $20\leq\lambda<1/$ 2, which shows that the smaller $\rho$, the smaller $(-\kappa_{1}/C_{D})$.
-Besides, all curves approach the same lower bound (purple curve) from the above in the limit $T\to-2$.
+1 (top) denotes the normal phase with $0\leq\lambda<1/2$, which shows that the smaller $\rho$, the smaller the $\kappa_{1}/C_{D}$.
+Besides, all curves approach the same lower bound (red line) from above in the limit $T\to-2$.
 On the other hand, Fig.
-1 (down) labels the singular phase with $1/2\leq\lambda\leq 1$.
-It shows that ($-\kappa_{1}/C_{D}$) with $\lambda>1/2$ is always larger than that of Einstein gravity (blue curve).
-Besides, the minimum of ($-\kappa_{1}/C_{D}$) increases with $\lambda$ and approaches zero when $\lambda\to 1$.
-Since we are interested in the lower bound of ($-\kappa_{1}/C_{D}$), we focus on the normal phase below.
-Figure 1: $(-\kappa_{1}/C_{D})$ in normal phase with $20\leq\lambda<1/$ 2 (up) and singular phase with $1/2<\lambda\leq 1$ (down).
-The up figure shows all curves approach the holographic limit (purple curve) from the above for $\rho\to-\infty$ ($T\to-2$) in the normal phase.
-The down figure shows $(-\kappa_{1}/C_{D})$ increases with $\lambda$ for $11/2<\lambda<$ 1.
-In the limit $\lambda\to 1$, we have $\kappa_{1}/C_{D}\to 0$ in the singular phase.
-$\lambda=1/2$ is the phase transition point.
+1 (bottom) labels the singular phase with $1/2\leq\lambda\leq 1$.
+It shows that $\kappa_{1}/C_{D}$ with $\lambda>1/2$ is always larger than that of Einstein gravity (blue curve).
+Besides, the minimum of $\kappa_{1}/C_{D}$ increases with $\lambda$ and approaches zero when $\lambda\to 1$.
+Since we are interested in the lower bound of ($\kappa_{1}/C_{D}$), we focus on the normal phase below.
+Let us explain the physical reasons why the minimal brane tension yields the lower bound of the Casimir effect in the normal phase.
+First, the Casimir energy equals the free energy at zero temperature, which can be calculated holographically by the Euclidean gravitational action.
+Naturally, the smaller the brane tension (energy), the smaller the Euclidean action, and, thus, the smaller the Casimir energy.
+Second, the brane tension decreases under boundary RG flow, i.e., $T_{\text{IR}}<T_{\text{UV}}$ Fujita:2011fp .
+At IR, many modes become inactive, resulting in reduced energy.
+Third, the Casimir effect is driven by long-range quantum correlations Dantchev:2022hvy .
+As the tension (energy) decreases, these quantum correlations become stronger, which in turn increases the Casimir amplitude and reduces the Casimir energy.
+In summary, the minimal brane tension corresponds to the minimal Casimir energy.
+Figure 1: $\kappa_{1}/C_{D}$ in normal phase with $0\leq\lambda<1/2$ (top) and singular phase with $1/2<\lambda\leq 1$ (bottom) for DGP gravity.
+Einstein gravity corresponds to $\lambda=0$.
+The top figure shows all holographic curves approach the same limit (red line) from above for $\rho\to$-$\infty$ ($T\to-2$) in the normal phase.
+The bottom figure shows $\kappa_{1}/C_{D}$ increases with $\lambda$ and approaches zero for $\lambda\to 1$ in the singular phase.
+Together, it shows all examples obey the holographic bound eq.(3) (red line).
 Now, we give some analytical discussions. Performing coordinate transformation $z=z_{\text{max}}y$ and expanding (11) around $x=\text{sech}^{2}(\rho)\to 0$, we derive perturbatively for normal phase
-$L_{I}=\frac{2\sqrt{\pi}\Gamma\left(\frac{4}{3}\right)}{\Gamma\left(\frac{5}{6}\right)}\Big{(}\frac{(1-2\lambda)^{2}}{x}\Big{)}^{\frac{1}{6}}+O\left(\frac{x}{(1-2\lambda)^{2}}\right)^{\frac{5}{6}}$ (9)
+$L_{\text{I}}=\frac{2\sqrt{\pi}\Gamma\left(\frac{4}{3}\right)}{\Gamma\left(\frac{5}{6}\right)}\Big(\frac{(1-2\lambda)^{2}}{x}\Big)^{\frac{1}{6}}+O\left(\frac{x}{(1-2\lambda)^{2}}\right)^{\frac{5}{6}}.$ (12)
 Taking limit $x\to 0$ ($T\to-2$) with $\lambda<1/2$, we obtain a universal limit independent of $\lambda$
-$\lim_{T\to-2}(\frac{-\kappa_{1}}{C_{D}})=\lim_{x\to 0}(\frac{-L_{\text{I}}^{3}}{C_{D}})=-\frac{\pi^{5/2}\Gamma\left(\frac{1}{3}\right)^{3}}{108\Gamma\left(\frac{5}{6}\right)^{3}}.$ (10)
+$\lim_{T\to-2}\frac{-\kappa_{1}}{C_{D}}=\lim_{x\to 0}\frac{-L_{\text{I}}^{3}}{C_{D}}=-\frac{\pi^{5/2}\Gamma\left(\frac{1}{3}\right)^{3}}{108\Gamma\left(\frac{5}{6}\right)^{3}}.$ (13)
 It verifies our claim that different holographic models give the same lower bound (3) in the normal phase.
-Note that (9) implies $L/\beta\to\infty$ for $x\to 0$ ($T\to-2$).
-Note also that the strip width $L\to Lz_{h}$ can remain finite when recovering $z_{h}\to 0$ in $h(z)$ (6).
-At the phase transition point $\lambda=1/2$, the lower bound (10) can be achieved by $L=\beta$.
-See appendix for derivations.
 The lower bound of $(-\kappa_{1}/C_{D})$ is saturated by negative brane tension.
-We remark that the negative brane tension is well-defined in AdS/BCFT Fujita:2011fp ; Takayanagi:2011zk ; Miyaji:2021ktr ; Tadashi .
-First, the brane tension is a cosmological constant rather than a kinetic-energy term on the end-of-the-wolrd (EOW) brane.
-Of course, the negative cosmological constant is well-defined in AdS/CFT.
+We remark that the negative brane tension is well defined in AdS/BCFT Fujita:2011fp ; Takayanagi:2011zk ; Miyaji:2021ktr ; Tadashi .
+First, the brane tension is a cosmological constant rather than a kinetic-energy term on the end-of-the-world brane.
+Of course, the negative cosmological constant is well defined in AdS/CFT.
 Second, negative brane tension generally yields a negative A-type boundary central charge.
@@
 Recall that the free scalar with the DBC has a negative A-type boundary central charge Jensen:2015swa .
-Tests — Let us test the holographic lower bound of the Casimir effect (3) by field theories.
-First, consider free CFTs.
-We have $(-\kappa_{1}/C_{D})=-3\pi\zeta(3)/8\approx-1.42$ for free fermions, and $(-\kappa_{1}/C_{D})=-\pi\zeta(3)/2\approx-1.89$ for free scalars with RBC and DBC.
-Next, consider CFTs with interactions.
-For Ising model with $\phi^{4}$ interaction at the critical point, we have $2\kappa_{1}\approx 0.820$ and $C_{D}\approx 0.193$ Toldin:2021kun ; Ising referee1 ; Ising referee2 , which yields $(-\kappa_{1}/C_{D})\approx-2.12$.
-Near the strip boundary, the scalar operator behaves as Cardy:1990xm
-$\langle\sigma(z)\rangle=\frac{a_{\sigma}}{(2z)^{\Delta_{\sigma}}}\Big{(}1+B_{\sigma}(\frac{z}{L})^{d}+...\Big{)},$ (11)
-where $\Delta_{\sigma}$ is the conformal dimension and $B_{\sigma}$ obeys the universal relation
-$\frac{B_{\sigma}}{\Delta_{\sigma}}=2^{d-1}(d-1)\pi^{-\frac{d}{2}}\Gamma\left(\frac{d}{2}\right)\ \frac{\kappa_{1}}{C_{D}}.$ (12)
-Here, we use the notation of Toldin:2021kun with $d=3$ for 3d CFTs.
-For $O(N)$ model with $N=2,3$, Toldin:2021kun obtains $B_{\sigma}\approx 1.21,1.07$ and $\Delta_{\sigma}\approx 0.519088,0.518920$ by applying Monte Carlo simulations.
-Then (12) gives $(-\kappa_{1}/C_{D})\approx-1.83,-1.62$ for $N=2,3$, respectively.
-We summarize all the above results in Table.1, which all obey the holographic bound (3).
-It strongly supports our proposal.
-Interestingly, (12) suggests the bound of the Casimir effect imposes an upper bound of $B_{\sigma}/\Delta_{\sigma}$.
-Below we briefly comment on various generalizations of our proposal (3).
-We show only key points and leave detailed discussions and open questions to future work.
-General boundaries — Holography is expected to set a bound of Casimir effect for general boundary shapes, not be limited to strips.
+General boundaries — Holography is expected to set a bound on the Casimir effect for general boundary shapes, not limited to strips.
 Let us provide some evidence.
@@
 On the other hand, we have $f(\Omega)\to(C_{D}\pi/64)(\pi-\Omega)$ in the smooth limit $\Omega\to\pi$ Miao:2024ddp .
-These two limits suggest
-$(\frac{-f(\Omega)}{C_{D}})\geq(\frac{-f(\Omega)}{C_{D}})_{\text{holo}},\ \text{for}\ 0<\Omega\leq\pi,$ (14)
-where the holographic bound is derived in the limit $T\to-2$.
-By applying the method of Miao:2024ddp and this letter, we numerically verify (14) holds.
+These two limits suggest AdS/BCFT with minimal brane tension sets the lower bound of the wedge Casimir effect
+$\frac{-f(\Omega)}{C_{D}}\geq\lim_{T\to-2}(\frac{-f(\Omega)}{C_{D}})_{\text{holo}},\ \text{for}\ 0<\Omega\leq\pi.$ (15)
+See the following long paper Miao:2025utb for the derivation and test of (15).
 Let us go on to discuss the general boundary.
-Near any smooth boundary, the normal-normal component of stress tensor takes a universal form in flat space Deutsch:1978sc ; Miao:2017aba
+Near any smooth boundary, the normal-normal component of the stress tensor takes a universal form in flat space Deutsch:1978sc ; Miao:2017aba
 $\langle T_{nn}(z)\rangle=\alpha\frac{\text{Tr}\bar{k}^{2}}{z}-\frac{\alpha}{2}k\text{Tr}\bar{k}^{2}\log(\frac{z}{L})+t_{nn}+O(z)$ (16)
@@
 A weaker proposal is for the average pressure
-$(\frac{\bar{t}_{nn}}{C_{D}})\geq(\frac{\bar{t}_{nn}}{C_{D}})_{\text{holo}},$ (16)
+$(\frac{\bar{t}_{nn}}{C_{D}})\geq\lim_{T\to-2}(\frac{\bar{t}_{nn}}{C_{D}})_{\text{holo}},$ (17)
+where $\bar{t}_{nn}=(\int_{\partial M}t_{nn}dS)/(\int_{\partial M}dS)$ is related to the variation of the total energy concerning the system size, i.e., dE/dL-dE/dL.
+General dimensions — Let us comment on the holographic bound of the Casimir effect for a strip in general dimensions.
+The corresponding Casimir amplitude $\kappa_{1}$ and norm of displacement operator $C_{D}$ are given by Miao:2024ddp for Einstein gravity.
+By taking the limit $T\to-(d-1)$, we derive a universal lower bound,
+$\frac{-\kappa_{1}}{C_{D}}\geq\frac{-2^{d-2}d\pi^{d-\frac{1}{2}}\Gamma\left(\frac{d-1}{2}\right)\Gamma\left(\frac{1}{d}\right)^{d}}{\Gamma(d+2)\left(d\Gamma\left(\frac{1}{2}+\frac{1}{d}\right)\right)^{d}}.$ (18)
 […]
-Higher dimensions — Let us comment on the holographic bound of Casimir effect for a strip in higher dimensions.
-The corresponding Casimir amplitude $\kappa_{1}$ and norm of displacement operator $C_{D}$ is given by Miao:2024ddp for Einstein gravity.
-By taking the limit $T\to-(d-1)$, we derive a universal lower bound
-$(\frac{-\kappa_{1}}{C_{D}})\geq\frac{-2^{d-2}d\pi^{d-\frac{1}{2}}\Gamma\left(\frac{d-1}{2}\right)\Gamma\left(\frac{1}{d}\right)^{d}}{\Gamma(d+2)\left(d\Gamma\left(\frac{1}{2}+\frac{1}{d}\right)\right)^{d}}.$ (17)
-[…]
+$(\frac{-\kappa_{1}}{C_{D}})|_{d=2}\geq-\frac{\pi^{3}}{12}.$ (19)
+According to Bloete:1986qm , the inequality is saturated for general 2D CFTs with the same boundary conditions on the two planes.
+On the other hand, if we impose different boundary conditions on the two planes, we get larger Casimir energy Bachas:2006ti ; Diatlyk:2024qpr
+$(-\kappa_{1})_{\text{differ BC}}\geq(-\kappa_{1})_{\text{same BC}}.$ (20)
+Note that $C_{D}=c/(2\pi^{2})$ is independent of the choices of boundary conditions, where $c$ is the central charge of 2d CFTs.
+Considering all the above statements, we get solid proof of the bound (19) for 2D CFTs, which is a strong support for our proposal (18).
 Non-CFTs — Interestingly, the lower bound of (3) also works for a general class of non-CFTs.
@@
 By applying the textbook method, we obtain for both RBC and DBC, i.e., $16\pi\kappa_{1}=-2M^{2}\log\left(1-e^{-2M}\right)+2M\text{Li}_{2}\left(e^{-2M}\right)+\text{Li}_{3}\left(e^{-2M}\right)$, where $M=mL>0$ is a dimensionless parameter, and Li denotes the polylogarithmic function.
-Remarkably, $\kappa_{1}$ is independent of $\xi$ and decreases with mass
+Remarkably, $\kappa_{1}$ is independent of $\xi$ and decreases with mass,
 $\frac{d\kappa_{1}}{dM}=\frac{M^{2}(1-\coth(M))}{8\pi}<0.$ (21)
-It means the Casimir effect is suppressed by mass Bordag:2009zz .
-For non-CFTs, we can define $C_{D}$ via $\langle D(y)D(0)\rangle=C_{D}/|y|^{6}+O(1/|y|^{5})$ in half space, where $D=T_{nn}(0)$ Billo:2016cpy .
-At small distance $m|y|\ll 1$, the mass effect can be ignored.
+For free non-CFTs, we can define $C_{D}$ via $\langle D(y)D(0)\rangle=C_{D}/|y|^{6}+O(1/|y|^{5})$ in half-space, where $D=T_{nn}(0)$ Billo:2016cpy .
+At a small distance $m|y|\ll 1$, the mass effect can be ignored.
 As a result, $C_{D}$ is independent of mass.
-The above discussions show $(-\kappa_{1}/C_{D})$ is larger than that of massless scalar, thus obeys the holographic lower bound (3).
-The above arguments apply to general massive free field theories.
-Discussions — In conclusion, we propose that holography sets a fundamental bound for the Casimir amplitude ratio to the norm of displacement operator.
-The Casimir amplitude is generally defined as the average pressure $\bar{t}_{nn}$ on the boundary.
-We take 3d CFTs in a strip as an example and then briefly discuss the generalization to other boundary shapes, higher dimensions, and non-CFTs.
+The above discussions show $\kappa_{1}/C_{D}$ is larger than that of a massless scalar, thus obeys the holographic lower bound (3).
+The above arguments apply to general massive free theories.
+The Casimir effect arises from long-range quantum correlations Dantchev:2022hvy .
+When mass is introduced, these correlations become short range, which generally reduces the Casimir amplitude $\kappa_{1}$ Bordag:2009zz .
+Recall that $C_{D}$ is mass independent.
+Consequently, $\kappa_{1}/C_{D}$ reaches its minimum value in the massless limit for general free theories.
+Discussions — In conclusion, we propose that holography sets a fundamental bound on the ratio of the Casimir amplitude to the displacement operator norm.
+The Casimir amplitude is defined as the average pressure $\bar{t}_{nn}$ on the boundary.
+We take 3D CFTs in a strip as an example and then briefly discuss the generalization to other boundary shapes, higher dimensions, and free non-CFTs.
 It is interesting to test our proposal in experiments of critical systems Krech:Casimir Effect .
-Finding a field-theoretical proof or counterexample for the holographic bound of the Casimir effect is also enjoyable.
-Usually, a bound is set by either free or strongly coupled theory.
-For entanglement entropy (EE), there is a similar bound set by free theory, i.e., $(-F_{0}/C_{T})\geq(-F_{0}/C_{T})_{\text{free scalar}}$ Bueno:2023gey , where $F_{0}$ is the universal coefficient of EE and $C_{T}$ is a bulk central charge.
-We remark that the dual brane for EE is tensionless while the one for the Casimir effect is tensive.
-With a tensive brane, we have more parameters to get a smaller lower bound of Casimir effect in holography than in free theory.
-Ising model of Table.
-1 also shows a free scalar does not set the lower bound of Casimir effect.
-In AdS3, conformal bootstrap restricts brane tension $|T|\leq 0.99$ sightly stronger than $|T|\leq 1$ Collier:2021ngi .
+Finding a field-theoretical proof or counterexample for our proposal is also enjoyable.
+Our findings help estimate how much the Casimir effect contributes to dark energy and whether sufficient Casimir energy can be generated to support a traversable wormhole macroscopically.
+Additionally, it helps evaluate the maximum Casimir force between nanodevices, which can enhance nanotechnology designs.
+Usually, a bound is set by either strongly coupled or free theories Kovtun:2004de ; Hofman:2008ar ; Hofman:2016awc .
+For instance, the KSS bound Kovtun:2004de for fluid is set by the strongly coupled theory dual to gravity, while the Hofman-Maldacena bound Hofman:2008ar ; Hofman:2016awc for bulk central charges is set by free theories.
+For entanglement entropy, there is a similar bound set by free theories, i.e., $(-F_{0}/C_{T})\geq(-F_{0}/C_{T})_{\text{free scalar}}$ Bueno:2023gey , where $F_{0}$ is the universal coefficient of entanglement entropy and $C_{T}$ is a bulk central charge.
+We remark that the dual brane for entanglement entropy is tensionless while the one for the Casimir effect is tensive.
+With a tensive brane, we have more parameters to get a smaller lower bound of the Casimir effect in holography than in free theory.
+The Ising model of Table 1 also shows that a free scalar does not set the lower bound of the Casimir effect.
+In AdS3, a conformal bootstrap restricts brane tension $|T|\leq 0.99$ sightly stronger than $|T|\leq 1$ Collier:2021ngi .
 If this is also the case for both $T$ and DGP coupling $\lambda$ in higher dimensions, it may give a slightly stronger lower bound of the Casimir effect.
 Anyway, this Letter takes a significant step toward exploring the lower bound of the Casimir effect.
-Many open problems remain to be explored.
+Many open problems remain to be explored, particularly the generalization to non-CFTs with interactions.
 Acknowledgements
 We thank T.
-Takayanagi, J.
+Takayanagi and J.
 X.
 Lu for valuable comments and discussions.
-We are grateful to the conference “Gauge Gravity Duality 2024”, where the work was completed.
+We are grateful to the conference “Gauge Gravity Duality 2024,”where the work was completed.
 This work is supported by the National Natural Science Foundation of China (No.12275366).
 Appendix A Bound of Casimir effect from DGP gravity
-This appendix studies the holographic bound of Casimir effect in DGP gravity. In particular, we give detailed derivations of eqs.(8, 10) in the main text.
-Let us quickly recall some key points. The vacuum of a strip is dual to AdS soliton
+This supplemental material studies the holographic bound of the Casimir effect in DGP gravity. We give detailed derivations of equations (11, 12, 13) in the main text.
+Let us quickly recall some key points. The vacuum of a strip is dual to the AdS soliton
 $ds^{2}=\frac{\frac{dz^{2}}{h(z)}+h(z)d\theta^{2}-dt^{2}+dy^{2}}{z^{2}},$ (22)
-where $h(z)=1-z^{3}/z_{h}^{3}$, and $\theta$ has the range $0\leq\theta\leq L$ on the AdS boundary $z=0$. We label the embedding function of brane $Q$ as
+where $h(z)=1-z^{3}/z_{h}^{3}$, and $\theta$ has the range $0\leq\theta\leq L$ on the AdS boundary $z=0$. We label the embedding function of the brane $Q$ as
 $\theta=S(z).$ (23)
-It obeies NBC
+It obeys NBC
 $K^{ij}-(K-T+\lambda\mathcal{R})h^{ij}+2\lambda\mathcal{R}^{ij}=0,$ (24)
-where the tension is parameterized as Miao:2023mui
+where the tension is parameterized as
 $T=2\tanh(\rho)-2\lambda\text{sech}^{2}(\rho).$ (25)
-This parameterization can be derived from the gravity dual of half space.
-Since the region near any smooth boundary can be approximately by half space at leading order of $z$, it works for general boundaries.
+This parameterization can be derived from the gravity dual of a half-space.
+Since the region near any smooth boundary can be approximated by a half-space at the leading order of $z$, it works for general boundaries.
 Substituting (23) into NBC (24), we derive near the two boundaries of strip
 $\theta=-\sinh(\rho)z+O(z^{2}),\ \ \theta-L=\sinh(\rho)z+O(z^{2}).$ (26)
-Let us first study the case with negative brane tension $T\leq 0$, the other case with $T>0$ can be obtained by the complement of the first case.
-This is the typical trick used in Fujita:2011fp .
+Let us first study the case with negative brane tension $T\leq 0$; the other case with $T>0$ can be obtained by the complement of the first case.
+It is the typical trick used in Fujita:2011fp .
 Let us explain more.
 See Fig.
-2 for the geometry of holographic dual of strip.
-The region between red/blue curves (branes) and black line is the bulk dual of strip I with $T\leq 0$; its complement in bulk is the gravity dual for strip II (green line) with $T>0$.
+2 for the geometry of the holographic dual of the strip.
+The region between red/blue curves (branes) and the black line is the bulk dual of strip I with $T\leq 0$; its complement in the bulk is the gravity dual for strip II (green line) with $T>0$.
 As shown in Fig.
-2, the gravity duals of strip I and strip II shares the same EOW branes (blue or red curve depending on the theory parameter).
+2, the gravity duals of strip I and strip II share the same EOW branes (blue or red curve depending on the theory parameter).
 As a result, the NBCs (24) for strip I and strip II take the same values.
 However, the extrinsic curvatures $K_{ij}$ flip signs, while the induced metric $h_{ij}$ and intrinsic Ricci tensor $\mathcal{R}_{ij}$ remain invariant when crossing the branes.
-In the viewpoint of NBC (21), $T$ and $\lambda$ equivalently change signs when crossing the branes.
+From the viewpoint of NBC (24), $T$ and $\lambda$ equivalently change signs when crossing the branes.
 Note that the bulk dual for strip II contains the ‘horizon’ $z=z_{h}$.
 To remove the conical singularity at $z=z_{h}$, we fix the period of angle $\theta$ as $\beta=4\pi/|h^{\prime}(z_{h})|=4\pi z_{h}/3$.
-Thus for $T>0$, the left and right vertical dotted lines of Fig.
+Thus, for $T>0$, the left and right vertical dotted lines of Fig.
 2 are identified due to the periodicity of angle $\theta$.
@@
 2 correspond to $\rho<0$ and $\rho>0$, respectively.
-We focus on the left half of branes, the right half can be obtained by symmetry.
+We focus on the left half of the branes; the right half can be obtained by symmetry.
 NBC (24) yields an independent equation for $T\leq 0$
 $T=-\frac{2h(z)\left(h(z)S^{\prime}(z)\sqrt{h(z)S^{\prime}(z)^{2}+\frac{1}{h(z)}}+\lambda\right)}{h(z)^{2}S^{\prime}(z)^{2}+1}.$ (27)
-Let us verify that it corresponds to the left half of branes.
-From (22,24), we derive $S^{\prime}(0)=-\sinh(\rho)$ near $\theta_{1}$ of Fig.
+Let us verify that it corresponds to the left half of the branes.
+From (25,27), we derive $S^{\prime}(0)=-\sinh(\rho)$ near $\theta_{1}$ of Fig.
 2, which agrees with the first equation of (26).
@@
 The negative sign of (29) corresponds to the right halves of branes.
-From (20,26), we derive the width of strip I for $T\leq 0$
+From (23,29), we derive the width of strip I for $T\leq 0$
 $L_{\text{I}}=\int_{0}^{z_{\max}}\frac{2\Big(\sqrt{h\left(z_{\max}\right)}-\lambda h(z)\Big)\ dz}{h(z)\sqrt{H(z,z_{\max})-h(z_{\max})}},$ (30)
 where $z_{\max}=(1-T^{2}/4)^{1/3}$ from (28).
-We stress that the above formula works for both the cases of blue curve and red curve of Fig.
+We stress that the above formula works for both the cases of the blue curve and the red curve of Fig.
 2.
 In particular, the integrand $2S^{\prime}(z)\sim 2\Big(\sqrt{h\left(z_{\max}\right)}-\lambda h(z)\Big)$ flips signs at turning point $z_{c}$ (28), which is the expected feature for the red curve.
-It suggests the integrand cannnot be chosen as the absolute value $2|S^{\prime}(z)|\sim 2\Big{|}\sqrt{h\left(z_{\max}\right)}-\lambda h(z)\Big{|}$, which would miss the case of red curve.
+It suggests the integrand cannot be chosen as the absolute value $2|S^{\prime}(z)|\sim 2\Big|\sqrt{h\left(z_{\max}\right)}-\lambda h(z)\Big|$, which would miss the case of the red curve.
 Figure 2: Geometry of holographic strip: a portion of AdS soliton.
@@
 $L=\begin{cases}\ L_{\text{I}},&\ \text{for}\ T\leq 0,\\ \ L_{\text{II}},&\ \text{for}\ T\geq 0.\end{cases}$ (32)
-One can check $L$ is continuous at $T=0$, which is a test of our results.
+One can check that $L$ is continuous at $T=0$, which is a test of our results.
 See Fig.
 3 for example.
-As mentioned in main text, the constraint $0\leq L$ set an upper bound of $\lambda$.
-It is clear from (27) that large $\lambda$ would give negative $L$.
+As mentioned in the main text, the constraint $0\leq L$ sets an upper bound of $\lambda$.
+It is clear from (30) that a large $\lambda$ would give a negative $L$.
 We numerically get the range $0\leq\lambda\leq 1$ for $L\geq 0$.
 As shown in Fig.
-3 , the case $20\leq\lambda\leq 1/$ 2 can continuously transform into that of Einstein gravity, while the case $1/2<\lambda\leq 1$ cannot.
+3, the case $0\leq\lambda\leq 1/2$ can continuously transform into that of Einstein gravity, while the case $1/2<\lambda\leq 1$ cannot.
 We name them normal and singular phases, respectively.
 $\lambda=1/2$ is the phase-transition point.
-Figure 3: Strip widths in normal phase with $20\leq\lambda<1/$ 2 (up) and in singular phase with $1/2<\lambda\leq 1$ (down).
+Figure 3: Strip widths in normal phase with $0\leq\lambda<1/2$ (top) and in singular phase with $1/2<\lambda\leq 1$ (bottom).
 Note that $L$ can be larger than $\beta$ for negative enough $\rho$ in the normal phase.
@@
 On the other hand, $L$ is always smaller than $\beta$ in the singular phase.
-In both cases, $0\leq L$ imposes an upper bound of $\rho$.
+$0\leq L$ imposes a lower bound of $\rho$ in the singular phase, and a upper bound of $\rho$ in both phases.
 Now we make some analytical discussions. Performing coordinate transformation $z=z_{\text{max}}y$ and expanding (30) around $x=\text{sech}^{2}(\rho)\to 0$, we derive perturbatively for normal phase
-[…]
+$L_{\text{I}}=\int_{0}^{1}dy\frac{2\sqrt[3]{1-2\lambda}}{\sqrt[6]{x}\sqrt{1-y^{3}}}+(1-2\lambda)\ O\left(\frac{x}{(1-2\lambda)^{2}}\right)^{\frac{5}{6}}$
 $=$ $\frac{2\sqrt{\pi}\Gamma\left(\frac{4}{3}\right)}{\Gamma\left(\frac{5}{6}\right)}\Big(\frac{(1-2\lambda)^{2}}{x}\Big)^{\frac{1}{6}}+(1-2\lambda)\ O\left(\frac{x}{(1-2\lambda)^{2}}\right)^{\frac{5}{6}}$
-The norm of displacement operator becomes for $\rho<0$
+The norm of the displacement operator becomes for $\rho<0$
 $C_{D}=\frac{32}{\frac{2\pi\lambda\sqrt{x}}{1-2\lambda\sqrt{1-x}}+\pi\cot^{-1}\left(\sqrt{\frac{1}{x}-1}\right)}.$ (34)
@@
 $L_{\text{I}}\to\frac{2\sqrt{\pi}\Gamma\left(\frac{4}{3}\right)}{\Gamma\left(\frac{5}{6}\right)}\Big(\frac{(1-2\lambda)^{2}}{x}\Big)^{\frac{1}{6}}.$ (36)
-Interestingly, the second term of (A) vanishes for $\lambda\to 1/2$. In fact, this is also the case for higher-order expansions. From (33) and $L_{\text{I}}=\beta=4\pi/3$, we solve
+Interestingly, the second term of (A) vanishes for $\lambda\to 1/2$. It is also the case for higher-order expansions. From (36) and $L_{\text{I}}=\beta=4\pi/3$, we solve
 $x=\frac{729\Gamma\left(\frac{4}{3}\right)^{6}}{64\pi^{3}\Gamma\left(\frac{5}{6}\right)^{6}}(1-2\lambda)^{2}+O(1-2\lambda)^{3}.$ (37)